# A YTD correspondence for constant scalar curvature metrics

### v1

@@ 1 The main results
The usefulness of this inequality is similar to that of the ‘Archimedean’ counterpart $K^{\beta}\leq K$, already discussed. Sending $\beta\to\infty$, and using Theorem 1.2, we obtain that $K\{u_{t}\}\leq K^{NA}(u)$. Since the reverse inequality is already known [61, Theorem 1.2], we obtain an alternative argument for the recently announced YTD theorem of Boucksom–Jonsson:
Theorem 1.4 (=Corollary 9.4).
[23] For any $u\in\mathcal{E}^{1,NA}$, one has
$K\{u_{t}\}=K^{NA}(u).$
Moreover, a unique csck metric exists in $c_{1}(L)$ if and only if there exists $\gamma>0$ such that
$K^{NA}(u)\geq\gamma J^{NA}(u),\ \ u\in\mathcal{H}^{NA}_{\mathcal{M}}.$
This result, together with its more general version allowing holomorphic vector fields obtained in [23], verifies a conjecture of C. Li who proved the existence direction [61].
This YTD theorem and our Theorem 1.1 imply that our notion of uniform $K^{\beta}$-stability is equivalent to uniform $K$-stability for model filtrations, as they are both equivalent to the existence of a unique cscK metric. Of independent interest, at the end of Section 9 we give an alternative direct proof of this equivalence that bypasses cscK metrics.
In light of the preceding discussion and the examples of [2], it is natural to wonder if these stability notions are equivalent to the traditional notion of uniform K-stability as suggested by Székelyhidi [81] and formulated by Dervan–Boucksom–Hisamoto–Jonsson [46, 17]:
Question 1.5.
Is uniform K-stability equivalent to uniform $K^{\beta}$-stability (for $\beta\gg 0$)?
By the inequality $K\geq K^{\beta}$ noted above, $K^{\beta}$-stability implies $K$-stability, mirroring that Ding stability implies $K$-stability in the Fano case [5]. The converse direction remains a subject for further investigation.
@@
To the best of our knowledge, our $K^{\beta}$ functionals have not been previously considered in the K-stability literature. We note, however, that [29, Section 4.1] considered a similar-flavored functional in their alternative proof of the convexity of the $K$ energy. Their construction involves an additional Banach-Saks averaging, together with a different type of quantization in the definition of their $\mathcal{M}_{k}$ functional. At present, it remains unclear whether variants of our $K^{\beta}$ functionals yield other new notions of stability, suggesting another intriguing direction for future investigation.
@@
Although we did not make explicit use of the expansion (5), it provided key motivation for our extensive reliance on transcendental quantization. Such expansions have already played an important role in influential studies of Kähler–Einstein type metrics by Wu and Jeffres-Mazzeo-Rubinstein [89], [59, Section 9] (cf. [77, (6.3)]), where they were employed to initiate a Ricci continuity path from -$\infty$, following a suggestion of Tian–Yau [87].
Our results about the $\beta$-entropy $Ent^{\beta}(\cdot)$ hold for any general compact Kähler manifold, but to stay brief, in this paper we do not study the relationship of this functional with holomorphic vector fields of the manifold. Instead, we refer the reader to [61, 57, 58, 23] for exciting recent developments in this direction.
@@ Acknowledgments.
K.Z. is especially grateful to J. Chu for discussions on the expansion of $u^{\beta}$ back in 2021. We are grateful to P. Mesquita Piccione, J. Han, M. Xia for answering numerous questions, and we thank Y. Liu and Y. Odaka for suggestions that improved the presentation.
@@
During the final stage of this project, we became aware of the related work announced by Boucksom–Jonsson [19]. Until then we had not considered model filtrations, but afterwards we naturally wondered about the relationships between their work and ours, resulting in the brief §9, where we give an alternative argument for their Theorem 1.4 using $K^{\beta}$ functionals. We are grateful to them for sharing their preprint [23] and also for related discussions. As far as we can tell, our approaches are fundamentally different. The work of Boucksom–Jonsson is a culmination of the non-Archimedean approach to K-stability, developed over the past decade (see [17, 10, 20, 21] etc.), and relies on an approximation argument at the level of model filtrations/non-Archimedean data, using [4, 48]. In contrast, our approach is ‘Archimedean’. We introduce novel $K^{\beta}$ functionals that approximate the $K$ energy directly via transcendental quantization at the level of potentials (inspired by [6]).
Paralleling our transcendental YTD correspondence Theorem 6.2, we acknowledge an independent work by Mesquita–Piccione and Witt Nyström [66], proving a different kind of transcendental YTD existence theorem, more in the spirit of [61], obtained through the resolution of non-Archimedean Monge–Ampère equations.
@@ 2 Preliminaries
$\mathcal{H}_{\omega}:=\{u\in C^{\infty}(X)\textup{ s.t. }\omega_{u}:=\omega+\mathrm{dd^{c}}u>0\},$
where $\mathrm{dd^{c}}:=\frac{i}{2\pi}\partial\bar{\partial}.$ The metric completion of $\mathcal{H}_{\omega}$ with respect to the $d_{1}$ metric [32] is denoted by $\mathcal{E}^{1}$, which is called the space of finite energy potentials [56]. More precisely, given $u,v\in\mathcal{E}^{1}$, the $d_{1}$-distance between $u$ and $v$ is given as
$d_{1}(u,v)=I(u)+I(v)-2I(P(u,v)),$
where $I(\cdot)$ denotes the Monge-Ampère energy (recall (3)) and $P(u,v)\in\mathcal{E}^{1}$ is the potential theoretic minimum of $u$ and $v$, i.e., the largest potential in $\mathcal{E}^{1}$ lying below $u$ and $v$. In particular, $d_{1}(0,u)=-I(u)$ when $u\leq 0$.
Given $a,b\in[-\infty,\infty]$ with $a<b$, a path $(a,b)\ni t\mapsto u_{t}\in\mathcal{E}^{1}$ is called a finite energy subgeodesic segment if $U_{z}(x):=u_{-\log|z|^{2}}(x)$ (with $x\in X$ and $z$ being a complex) satisfies
$p_{1}^{*}\omega+\mathrm{dd}^{c}_{X\times{\mathbb{C}}}U_{z}\geq 0$
as a current on $X\times\{e^{-a/2}<|z|<e^{-b/2}\}$, where $p_{1}:X\times{\mathbb{C}}\to X$ is the projection map.
Let $h_{0},h_{1}\in\mathcal{E}^{1}$. As shown in [32], a distinguished $d_{1}$-geodesic $0,1]\ni t\mapsto h_{t}\in\mathcal{E}^{1}_{\omega}$ connecting $h_{0},h_{1}$ can be obtained as Perron envelope of finite energy subgeodesics:
$h_{t}:=\sup\{v_{t}\ |\ [0,1]\ni l\ \mapsto v_{l}\in\mathcal{E}^{1}\textup{ is a subgeodesic with }\limsup_{l\to 0,1}v_{l}\leq h_{0,1}\},\ t\in[0,1].$ (13)
@@
A subgeodesic ray $\{u_{t}\}$ is called sublinear if $u_{t}\leq Ct$ for some $C>0$. Recall that a geodesic ray is automatically sublinear due to [40, Theorem 3.7(iv)]. Finally, following the formalism of [37], by $\mathcal{R}^{1}$ we denote the space of finite energy geodesic rays emanating from 0, which is complete with respect to the chordal metric $d_{1}^{c}$ defined in (10) [37, Theorem 1.4].
A geodesic ray $\{u_{t}\}\in\mathcal{R}^{1}$ is bounded if $u_{t}\in\mathcal{E}^{1}\cap L^{\infty},\ t\geq 0$. [33, Theorem 1] implies that such rays are automatically linearly bounded: there exists $C>0$ such that $|u_{t}|\leq Ct,\ t\geq 0$.
@@
Test configurations.
@@
In practice, one often encounters semi-ample test configurations, for which $\mathcal{L}$ is only assumed to be relatively semi-ample. By [17, §2.5], such a generalization is hardly essential. Indeed, semi-ample test configurations are simply pullbacks of their ample models, and both induce the same geodesic ray. As a result, in this work, slightly abusing precision, we think of ample and semi-ample test configurations as the same objects.
@@
Conversely (see [17, §2.6]), for any semi-ample test configuration $(\mathcal{X},\mathcal{L})$, after replacing $\mathcal{L}$ with $\mathcal{L}+a\mathcal{X}_{0}$ for appropriate $a\in\mathbb{Q}$, the resulting test configuration has the same Futaki invariant, but induces a flag ideal $\mathfrak{a}$ on $X\times{\mathbb{C}}$ such that the sheaf $p_{1}^{*}(mL)\otimes\mathfrak{a}$ is globally generated for some divisible $m>0$. This ideal gives rise to a semi-ample test configuration $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ with a morphism $\mathcal{X}^{\prime}\xrightarrow{\tau}\mathcal{X}$ such that $\tau^{*}\mathcal{L}=\mathcal{L}^{\prime}$.
Most importantly, to any semi-ample test configuration $\mathcal{T}$ one can attach a $C^{1,1}$ geodesic ray $\{u_{t}^{\mathcal{T}}\}$ [72, 73, 76, 5, 30]. When the semi-ample test configuration is obtained via a flag ideal $\mathfrak{a}$ as described above, one further sees that the singularity type of $U^{\mathcal{T}}_{z}:=u^{\mathcal{T}}_{-\log|z|^{2}}$ near $X\times\{0\}$ is comparable to $\mathfrak{a}^{1/m}$, in the sense that
@@
where $\{f_{i}\}$ is a finite set of local generators of $\mathfrak{a}$. For this reason, $\{u_{t}^{\mathcal{T}}\}$ is also called a geodesic ray with algebraic singularities in the terminology of [10, §4.4].
Demailly regularization.
@@
We recall an important construction going back to [44, 10]. Let $\{u_{t}\}$ be a finite energy subgeodesic ray with $u_{t}\leq 0$. Put $U:=U_{z}=u_{-\log|z|^{2}}$, which extends to a quasi-plurisubharmonic (qpsh) function on $X\times\Delta$. For any $m\geq 1$, let $\mathfrak{a}_{m}:=\mathcal{I}(mU)$ be the multiplier ideal sheaf of $mU$. Note that $\mathfrak{a}_{m}$ is trivial away from $X\times\{0\}$ (by [94]), so it extends to a flag ideal on $X\times{\mathbb{C}}$. As shown in [10, Lemma 5.6], one can find a sufficiently large $m_{0}>0$, depending only on $(X,L)$, such that $p_{1}^{*}((m+m_{0})L)\otimes\mathfrak{a}_{m}$ is globally generated for any $m\geq 1$. Let $\mathcal{X}_{m}\xrightarrow{\mu_{m}}X\times{\mathbb{C}}$ be the normalized blowup along $\mathfrak{a}_{m}$ with exceptional divisor $E_{m}$. Put $\mathcal{L}_{m}:=\mu_{m}^{*}p_{1}^{*}L-\frac{1}{m+m_{0}}E_{m}$. Then $\mathcal{T}_{m}:=(\mathcal{X}_{m},\mathcal{L}_{m})$ is a semi-ample test configurations, inducing a geodesic ray $\{u_{t}^{\mathcal{T}_{m}}\}$. Using (14), one then observes that
$\int_{X\times\Delta}e^{(m+m_{0})U^{\mathcal{T}_{m}}-mU}\omega^{n}\wedge idz\wedge d\bar{z}<\infty.$ (15)

### v2

@@ 1 The main results
The usefulness of this inequality is similar to that of the ‘Archimedean’ counterpart $K^{\beta}\leq K$, already discussed. Sending $\beta\nearrow\infty$ and using $K^{\beta}\{u_{t}\}\nearrow K\{u_{t}\}$ (Theorem 6.3), we obtain that $K\{u_{t}\}\leq K^{NA}(u)$. Since the reverse inequality is known [70, Theorem 1.2], we obtain an alternative argument for the following key formula of Boucksom–Jonsson [27]:
$K\{u_{t}\}=K^{NA}(u),\ \ u\in\mathcal{E}^{1,NA}.$ (17)
Coupled with the results of [70] and additional non-Archimedean techniques, this formula yields the following YTD correspondence, also obtained in [27]:
Theorem 1.3.
[27] Let $(X,L)$ be a polarized manifold. The following are equivalent:
(i) $(X,L)$ admits a cscK metric in $c_{1}(L)$.
(ii) ${\mathbb{G}}:=\mathrm{Aut}_{0}(X,L)$ is reductive and $(X,L)$ is $\mathbb{G}$}-uniform K-stable with respect to models.
For the definition of $\mathbb{G}$}-uniform K-stability (with respect to test configurations, models, or log discrepancy models), we refer to (89). This result confirmed a conjecture of C. Li, who previously proved the existence direction in [70]. In fact, [27] carries out a considerably more refined analysis, establishing YTD-type correspondences involving $K$-polystability and prescribed symmetry in the framework of weighted cscK metrics as well.
By employing the framework provided by $K^{\beta}$-stability, we will refine Theorem 1.3, demonstrating that the existence of cscK metrics can be tested using only the distinguished subclass of models introduced in (13) (see Theorem 1.6 below). This will be achieved by making progress on a well-known regularization conjecture.
As emphasized in a series of papers [23, 26, 25, 70, 71] of Boucksom–Jonsson and Chi Li, the remaining step for addressing the strong form of the uniform YTD conjecture – as formulated in [91, 50, 21, 70] – is to establish a regularization property of the radial entropy:
Conjecture 1.4 (Boucksom–Jonsson).
Let $(X,L)$ be a polarized manifold with $\omega\in c_{1}(L)$. Let $\{u_{t}\}\in\mathcal{R}^{1}$ be a geodesic ray with $Ent\{u_{t}\}<\infty$. Then there exists a sequence of geodesic rays $\{v_{t,k}\}$ arising from test configurations of $(X,L)$ such that
$\lim_{k\to\infty}d_{1}^{c}(\{u_{t}\},\{v_{t,k}\})=0\quad\text{and}\quad Ent\{u_{t}\}=\lim_{k\to\infty}Ent\{v_{t,k}\}.$
Regarding this conjecture, remarkable progress was made by Chi Li [70, Proposition 6.3], showing that one can carry out the approximation via a sequence of geodesic rays $\{v_{t,k}\}$ attached to models. Li constructs these approximating models implicitly, by solving non-Archimedean Calabi–Yau equations [19].
Using our $\beta$-entropy we refine Li’s approximation result and make the construction more explicit. Specifically, we construct a regularizing sequence of log discrepancy models obtained as small perturbations of the test configurations arising from the classical multiplier ideal sheaf approximation procedure of [9] (recalled in §2)).
Theorem 1.5 (Theorem 10.1).
Let $(X,L)$ be a polarized manifold with $\omega\in c_{1}(L)$. Let $\{u_{t}\}\in\mathcal{R}^{1}$ be a geodesic ray with $Ent\{u_{t}\}<\infty$. Then there exists a sequence of smooth test configurations $(\mathcal{X}^{(k)},\mathcal{L}^{(k)})$ dominating $X\times{\mathbb{P}}^{1}$ and rational numbers $\beta_{k}\to\infty$ such that the geodesic rays $\{v_{t,k}\}$ attached to the log discrepancy models $(\mathcal{X}^{(k)},\mathcal{L}^{(k)}_{\beta_{k}})$ satisfy
$\lim_{k\to\infty}d_{1}^{c}(\{u_{t}\},\{v_{t,k}\})=0\quad\text{and}\quad Ent\{u_{t}\}=\lim_{k\to\infty}Ent\{v_{t,k}\}.$
With this regularization theorem in hand, the techniques of [43] and [41] immediately imply the following $\mathbb{G}$}-uniform YTD correspondence, sharpening Theorem 1.3:
Theorem 1.6 (Theorem 10.4).
Let $(X,L)$ be a polarized Kähler manifold. The following are equivalent:
(i) $(X,L)$ admits a cscK metric in $c_{1}(L)$.
(ii) ${\mathbb{G}}:=\mathrm{Aut}_{0}(X,L)$ is reductive and $(X,L)$ is $\mathbb{G}$}-uniform K-stable with respect to log discrepancy models.
Remark 1.7.
Fix any $\beta_{0}\gg 1$. Our argument shows that for condition (ii), it is sufficient to consider log discrepancy models $(\mathcal{X},\mathcal{L}_{\beta})$ with $\beta>\beta_{0}$ (recall (13)). Thus, it is enough to work with models that are in some sense ‘arbitrarily close’ to being test configurations. That being said, it remains unclear to what extent this refinement could lead to verifiable criteria for stability in concrete situations, suggesting avenues for future research.
In light of the above theorems, to resolve the strong form of the ($\mathbb{G}$}-)uniform YTD conjecture with respect to test configurations (see, for instance, [70, Conjecture 2.26]), it remains to determine whether one can regularize geodesic rays $\{u_{t}\}$ associated with log discrepancy models in the sense of Conjecture 1.4. At first sight, this appears to be more realistic, but time will tell whether the concrete structure of these log-discrepancy models will play a decisive role in the resolution of Conjecture 1.4.
It would also be desirable to extend Theorem 1.6 to the transcendental setting. This seems to require more work particularly when $\mathbb{G}$} is non-trivial, since an equivariant non-Archimedean theory for Kähler manifolds has yet to be fully developed (cf. [45, 75, 76]).
@@
To the best of our knowledge, our $K^{\beta}$ functionals have not been previously considered in the K-stability literature. We note, however, that [33, Section 4.1] considered a similar-flavored functional in their alternative proof of the convexity of the $K$ energy. Their construction involves an additional Banach-Saks averaging, together with a different type of quantization in the definition of their $\mathcal{M}_{k}$ functional. At present, it remains unclear whether variants of our $K^{\beta}$ functionals yield other novel notions of stability.
@@
Although we did not make explicit use of the expansion (6), it provided key motivation for our extensive reliance on transcendental quantization. Such expansions have already played an important role in influential studies of Kähler–Einstein type metrics by Wu and Jeffres-Mazzeo-Rubinstein [99], [67, Section 9] (cf. [87, (6.3)]), where they were employed to initiate a Ricci continuity path from -$\infty$, following a suggestion of Tian–Yau [97].
In this work we did not study the more general extremal metrics and their weighted versions. Instead, we refer the reader to [64, 65, 27] for exciting recent developments in this direction.
@@ Acknowledgments.
K.Z. is especially grateful to J. Chu for discussions on the expansion of $u^{\beta}$ back in 2021. We are grateful to P. Mesquita Piccione, J. Han, M. Xia for answering numerous questions, and we thank Y. Liu, Y. Odaka, Y.A. Rubinstein for suggestions that improved the presentation.
@@
During the final stage of this project, we became aware of related work announced by Boucksom–Jonsson [27]. Until then we had not considered working with Chi Li’s models, or exploring connections with non-Archimedean geometry, but afterwards we naturally wondered about the relationships between their work and ours. This line of thought led to our alternative argument for their key formula (17) using $K^{\beta}$ energies, and the $\mathbb{G}$–uniform YTD correspondence involving log discrepancy models (Theorem 1.6), making contact with their Theorem 1.3. We are grateful to them for sharing their preprint [27] and also for related discussions. As far as we can tell, our approaches are fundamentally different. The work of Boucksom–Jonsson is a culmination of the non-Archimedean approach to K-stability, developed over the past decade (see [21, 9, 23, 25] etc.), and relies on an approximation argument at the level of model filtrations/non-Archimedean data, using [3, 53]. In contrast, our approach is largely ‘Archimedean’. We introduce novel $K^{\beta}$ energies that approximate the $K$ energy directly via transcendental quantization at the level of potentials, inspired by [5].
Paralleling our transcendental YTD correspondence Theorem 5.5, we acknowledge independent work by Mesquita–Piccione and Witt Nyström [76], proving a different kind of transcendental YTD existence theorem, more in the spirit of [70], obtained through the resolution of non-Archimedean Monge–Ampère equations.
@@ 2 Preliminaries
$\mathcal{H}_{\omega}:=\{u\in C^{\infty}(X,\mathbb{R})\textup{ s.t. }\omega_{u}:=\omega+\mathrm{dd^{c}}u>0\},$
where $\mathrm{dd^{c}}:=\frac{i}{2\pi}\partial\bar{\partial}.$ The metric completion of $\mathcal{H}_{\omega}$ with respect to the $d_{1}$ metric [36] is denoted by $\mathcal{E}^{1}$, which is called the space of finite energy potentials [63]. More precisely, given $u,v\in\mathcal{E}^{1}$, the $d_{1}$-distance between $u$ and $v$ is given as $d_{1}(u,v)=I(u)+I(v)-2I(P(u,v)),$ where $I(\cdot)$ denotes the Monge–Ampère energy (recall (3)) and $P(u,v)\in\mathcal{E}^{1}$ is the potential theoretic minimum of $u$ and $v$, i.e., the largest potential in $\mathcal{E}^{1}$ lying below $u$ and $v$. In particular, $d_{1}(0,u)=-I(u)$ when $u\leq 0$.
Given $a,b\in[-\infty,\infty]$ with $a<b$, a path $(a,b)\ni t\mapsto u_{t}\in\mathcal{E}^{1}$ is called a finite energy subgeodesic segment if $U_{z}(x):=u_{-\log|z|^{2}}(x)$ (with $x\in X$ and $z$ being a complex) satisfies $p_{1}^{*}\omega+\mathrm{dd}^{c}_{X\times{\mathbb{C}}}U_{z}\geq 0$ as a current on $X\times\{e^{-a/2}<|z|<e^{-b/2}\}$, where $p_{1}:X\times{\mathbb{C}}\to X$ is the projection map.
Let $h_{0},h_{1}\in\mathcal{E}^{1}$. As shown in [36], a distinguished $d_{1}$-geodesic $(0,1)\ni t\mapsto h_{t}\in\mathcal{E}^{1}_{\omega}$ connecting $h_{0},h_{1}$ can be obtained as Perron envelope of finite energy subgeodesics:
$h_{t}:=\sup\{v_{t}\ |\ (0,1)\ni l\ \mapsto v_{l}\in\mathcal{E}^{1}\textup{ is a subgeodesic with }\limsup_{l\to 0,1}v_{l}\leq h_{0,1}\},\ t\in(0,1).$ (18)
@@
A subgeodesic ray $\{u_{t}\}$ is called sublinear if $u_{t}\leq Ct$ for some $C>0$. Recall that a geodesic ray is automatically sublinear due to [44, Theorem 3.7(iv)]. Finally, following the formalism of [41], by $\mathcal{R}^{1}$ we denote the space of finite energy geodesic rays emanating from 0, which is complete with respect to the chordal metric $d_{1}^{c}$ defined in (11) [41, Theorem 1.4].
A geodesic ray $\{u_{t}\}\in\mathcal{R}^{1}$ is bounded if $u_{t}\in\mathcal{E}^{1}\cap L^{\infty},\ t\geq 0$. [37, Theorem 1] implies that such rays are automatically linearly bounded: there exists $C>0$ such that $|u_{t}|\leq Ct,\ t\geq 0$.
@@
Test configurations and geodesic rays.
@@
In practice, one often encounters semi-ample test configurations, for which $\mathcal{L}$ is only assumed to be relatively semi-ample. For brevity, we will refer to these simply as test configurations. By [21, §2.5], such a generalization is hardly essential. Indeed, a semi-ample test configuration is simply the pullback of an ample one, and they both induce the same geodesic ray. As a result, in this work, slightly abusing precision, we think of ample and semi-ample test configurations as the same objects, and just call them as test configurations.
@@
Conversely (see [21, §2.6]), for any test configuration $(\mathcal{X},\mathcal{L})$, after replacing $\mathcal{L}$ with $\mathcal{L}+a\mathcal{X}_{0}$ for appropriate $a\in\mathbb{Q}$ (where $\mathcal{X}_{0}\subset\mathcal{X}$ denotes the central fiber over $0\in{\mathbb{C}}$), the resulting test configuration has the same Futaki invariant, but induces a flag ideal $\mathfrak{a}$ on $X\times{\mathbb{C}}$ such that the sheaf $p_{1}^{*}(mL)\otimes\mathfrak{a}$ is globally generated for some divisible $m>0$. This ideal gives rise to a test configuration $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ with a morphism $\mathcal{X}^{\prime}\xrightarrow{\tau}\mathcal{X}$ such that $\tau^{*}\mathcal{L}=\mathcal{L}^{\prime}$.
Most importantly, to any test configuration $\mathcal{T}=(\mathcal{X},\mathcal{L})$ one can attach a $C^{1,1}$ geodesic ray $\{u_{t}^{\mathcal{T}}\}$ [82, 83, 86, 4, 34]. After adding a multiple $a\mathcal{X}_{0}$ to $\mathcal{L}$ (which corresponds to adding $at$ to $u_{t}^{\mathcal{T}}$), we can assume that $(\mathcal{X},\mathcal{L})$ is obtained via a flag ideal $\mathfrak{a}$ as described above. Then from the construction of $\{u_{t}^{\mathcal{T}}\}$ the singularity type of $U^{\mathcal{T}}_{z}:=u^{\mathcal{T}}_{-\log|z|^{2}}$ near $X\times\{0\}$ is comparable to $\mathfrak{a}^{1/m}$, in the sense that
@@
where $\{f_{i}\}$ is a finite set of local generators of $\mathfrak{a}$. For this reason, $\{u_{t}^{\mathcal{T}}\}$ is called a geodesic ray with algebraic singularities, in the terminology of [9, §4.4].
Compactification and domination.
In practice, it is also useful to compactify a test configuration $(\mathcal{X},\mathcal{L})$ trivially over $\infty\in{\mathbb{P}}^{1}$ (see [101, Definition 2.7]), which we still denote as $(\mathcal{X},\mathcal{L})$. In this paper we assume that all test configurations $(\mathcal{X},\mathcal{L})$ are compactified in this sense, making intersection theory more convenient on $\mathcal{X}$. In this context $\mathcal{L}$ being relatively (semi)ample means that $\mathcal{L}+c\mathcal{X}_{0}$ is (semi)ample for some $c\in\mathbb{Q}_{>0}$.
Most of our discussions will involve birational invariance, hence we will often assume that the total space $\mathcal{X}$ of a test configuration $(\mathcal{X},\mathcal{L})$ is smooth, where $\mathcal{X}_{0}$ has simple normal crossing support. This can always achieved after taking a resolution (which does not affect the associated geodesic ray). Such $(\mathcal{X},\mathcal{L})$ is called a smooth model (of the original test configuration).
[…]
Let $(\mathcal{X},\mathcal{L})$ be a smooth test configuration dominating $X\times{\mathbb{P}}^{1}$, say $\mathcal{X}\xrightarrow{\pi}X\times{\mathbb{P}}^{1}$. Denote by $p_{1}:X\times{\mathbb{P}}^{1}\to X$ the projection map. By [90, Proposition 3.10] there is a uniquely determined $\mathbb{Q}$-Cartier divisor $D$ supported in $\mathcal{X}_{0}$ such that
$\pi^{*}p_{1}^{*}L-\mathcal{L}\sim_{\mathbb{Q}}D.$ (20)
Let $\{u_{t}\}$ be the geodesic ray associated to $(\mathcal{X},\mathcal{L})$. By the analysis in [90, Lemma 4.6] and [34, §3] the singularity type of $U_{z}:=u_{-\log|z|^{2}}$ when pulled back to $\mathcal{X}$ is $L^{\infty}$-compatible with the divisor $D$, in the sense of [90, Definition 4.5] (up to a sign). When $D$ is effective, this means that we can find an $S^{1}$-invariant qpsh function $\psi_{D}$ in a neighborhood of $\mathcal{X}_{0}$ such that
$\theta_{D}+\mathrm{dd^{c}}\psi_{D}=[D],\text{ and }|\pi^{*}U_{z}-\psi_{D}|=O(1),$ (21)
where $\theta_{D}$ is a smooth $(1,1)$-form on $\mathcal{X}$ representing the cohomology class $\{D\}$.
Demailly regularization of rays.
We recall an important approximation procedure going back to [48, 9]. Let $\{u_{t}\}$ be a finite energy subgeodesic ray with $u_{t}\leq 0$. Put $U:=U_{z}=u_{-\log|z|^{2}}$, which extends to a quasi-plurisubharmonic (qpsh) function on $X\times\Delta$.
@@
For any $m\geq 1$, let $\mathfrak{a}_{m}:=\mathcal{I}(mU)$ be the multiplier ideal sheaf of $mU$. Note that $\mathfrak{a}_{m}$ is trivial away from $X\times\{0\}$ (by [105]), so it extends to a flag ideal on $X\times{\mathbb{C}}$. As shown in [9, Lemma 5.6], one can find a sufficiently large $m_{0}>0$, depending only on $(X,L)$, such that $p_{1}^{*}((m+m_{0})L)\otimes\mathfrak{a}_{m}$ is globally generated for any $m\geq 1$. Let $\mathcal{X}_{m}\xrightarrow{\mu_{m}}X\times{\mathbb{C}}$ be the normalized blowup along $\mathfrak{a}_{m}$ with exceptional divisor $E_{m}$. Put $\mathcal{L}_{m}:=\mu_{m}^{*}p_{1}^{*}L-\frac{1}{m+m_{0}}E_{m}$. Then $\mathcal{T}_{m}:=(\mathcal{X}_{m},\mathcal{L}_{m})$ is a test configuration of $(X,L)$, inducing a geodesic ray $\{u_{t}^{\mathcal{T}_{m}}\}$. Using (19), one then observes that
$\int_{X\times\Delta}e^{(m+m_{0})U_{z}^{\mathcal{T}_{m}}-mU_{z}}\omega^{n}\wedge idz\wedge d\bar{z}<\infty.$ (22)